# A High-Resolution, US-scale Digital Similar of Interacting Livestock, Wild Birds, and Human Ecosystems with Applications for Multi-host Epidemic Spread

### v2

@@ 2.1 The Digital Similar
[…]
@@
This component provides a grid-level representation of the human population with emphasis on agricultural workers (workers associated with livestock and its processing). For each cell $v$$\pi(v,\delta,\epsilon)$ denotes the size of the subpopulation that belongs to the demographic group $\delta\in\Delta$ defined by attributes including age group and sex and employed in professional classes specified by $\epsilon\in\mathcal{E}$. An employment group $\epsilon$ is defined by occupation and industry attributes, where non-agricultural employees are all binned into one group, namely non-agriculture.
@@ 2.3 Validation of farm locations
Concentrated Agricultural Feed Operations (CAFOs) are large animal feeding operations that are a potential hazard to the environment and health. CAFOs are regulated by the Environmental Protection Agency (EPA), and some state agencies provide location information, among other attributes. We obtained such data from CafoMaps [49]. We focus on cattle, hogs, and chickens. For each county–livestock instance for which such data is present, we selected large farms from our farm assignment based on livestock-specific thresholds informed by CAFO size specifications provided by various states [52]. We computed the Haversine distance of each CAFO location to the centroid of each grid cell that contains a large farm. We construct a weighted complete bipartite graph $G(A,B)$ for each county–livestock instance. Here $A$ corresponds to farm locations from our assignments; each farm is assumed to be located at the centroid of the grid cell to which it belongs. The set $B$ corresponds to CAFO locations. For each $u\in A$ and $v\in B$, the weight on the edge $(u,v)$ is the inverse of the distance between the two locations. We compute a maximum weighted perfect matching [^2] For a complete bipartite graph $G(A,B,E)$, where $|A|=m$ and $|B|=n$, a perfect matching consists of $\min\{m,n\}$ edges. of this bipartite graph to match each CAFO location to a farm in $\mathcal{DS}$. The main objective is to map as many CAFO locations as possible. It is possible that the number of farms is greater than the number of CAFO locations, as not all locations are listed. The results of the matching are analyzed in Figure 5(a). We considered two sets of thresholds, the second set corresponding to larger farms compared to the first. A large percentage of CAFO locations were matched in the case of cattle ($>95$%) and chickens ($>83$%), while in the case of hogs we observe only $50$% match. A closer examination of AgCensus data reveals the reason for the low number of matches for hogs: the number of farms specified by the AgCensus dataset for the relevant size categories is less than the number of CAFO locations specified. Among the matched locations, we observe that $90$% of the CAFO locations are at most $10$ miles from the grid centroid of the corresponding farm from the digital similar, which places it in the same grid cell or a neighboring one.
@@
We note that a majority of matched CAFO farms corresponding to cattle and hogs are within $20$ miles of the matched farm in $\mathcal{DS}$.
@@ 2.4 Wild bird abundance and H5N1 incidence
The results presented in Figure 5(b) summarizes the Top-K recall results for all states and different parts of the year. We observe that, for all quarters, the larger the bird population, the greater the likelihood of observed H5N1 cases. Secondly, we observe that, in most cases, the top 10 abundant birds cover all reported H5N1 cases (for $k\geq 10$), so the recall value is close to $1$. More plots are shown in the supplement for multiple US states with details about specific bird types. These results suggest a robust correlation between bird abundance and H5N1 case occurrences.
@@ 2.6 Risk Estimation
[…]
@@
Figure 7: Results corresponding to livestock subtype-specific quarterly risk maps based on county ranks. The risk scores were ranked per quarter and counties were assigned risk profiles: “Very high” ($\geq 95\%$), “High” ($90-94$), “Medium” ($75-89$) and “Low”. Here, we show summary plots for milk cattle and turkeys. (a) Persistence of elevated risk: Quarter-specific ranks were combined to obtain a single ranking of counties by sorting counties by number of occurrences of “Very high” risk profiles across quarters followed by number of occurrences of “high” risk profiles and so on. (b) Period of peak risk: For each county, we plot the period for which the risk is maximum ($\operatorname*{arg\,max}_{t={1,2,3,4}}R_{c}(s,t)$). (c) Comparing ground-truth incidence reports with risk profiles. Each incidence is mapped to the risk profile of the corresponding county–quarter pair. Then, the incidences are binned into the respective risk profiles. (d) Persistence of elevated risk under a potential scenario of H5N1 spread in cattle (definition same as in (a)).
@@ 4.2 Livestock
AgCensus provides counts of heads (i.e., population size) and farms for various livestock types and subtypes. The data is available at three different administrative levels – country, state, and county. The data organization is livestock-type specific, making it a non-trivial task to extract relevant information. Farms are binned into categories based on the head counts of the corresponding livestock type. We preprocessed the data to ensure that these categories are disjoint and the categorization is identical at both state and county levels. Ideally, given a livestock subtype and administrative level, both farm and head counts are provided for each farm size category. Also provided is the total count of farms and heads. Together, we have four possible types of counts: (i) state-total: total number of farms/heads; (ii) state-by-farm-size: number of farms/heads per farm size category, and corresponding county-level counts; (iii) county-total; and (iv) county-by-farm-size. More details with examples are provided in the supplement. However, some head counts are missing in all count types. In the case of poultry, even farm size categories are missing for all subtypes except for (chicken) layers. The GLW [24] dataset provides a gridded distribution of livestock abundance at $5$ arc minute resolution for the livestock types, but not for the subtypes. The details of the processing of the GLW data are provided in the supplement.
@@ 4.3 Wild Bird Abundance and Movement
We leveraged data from eBird’s Status and Trends products [27] (see eBird in Table 1) to construct this component. We recall that this component $\mathcal{B}\big{(}\Theta_{\mathcal{B}},A(\cdot),G_{\mathcal{B}}\big{)}$ captures spatiotemporal abundance and movement of multiple species of birds. The eBird data provides weekly estimates of relative bird abundance across a high-resolution grid (2.96km×$\times$2.96km). These estimates, derived from ensemble machine learning models, represent the expected count of a species on a standardized eBird checklist at a given location and date. The models combine millions of citizen science observations with environmental predictors, accounting for factors such as land cover, climate variables, and observation effort. A relative abundance of 1.0 for a species at a particular location and time would indicate that an average eBird checklist at that place and time would be expected to count one individual of that species. Higher values indicate more individuals would be expected, while lower values indicate the species would be observed less frequently or in smaller numbers. This approach ensures that our model captures the most relevant species for studying avian influenza transmission. We chose bird species for which H5N1 cases were observed in the period 2022-2024. A total of 40 species of birds were identified, of which abundance data was available for 36 species. For each of the selected species, we extract relative abundance values along with their associated geographic coordinates for all 52 weeks in the year. Our processing pipeline is described in the supplement.
@@ 4.5 Human Population
We develop a gridded representation of the US population with rich demographic and employment-related attributes. This data is derived from a synthetic population [18, 32, 23] that is developed using diverse datasets such as census data, land use data, activity patterns, building maps, etc. Each individual in the population is associated with a residential location, an occupation identified by the Standard Occupational Classification code (SOCP), and an industry, identified by the North American Industry Classification System (NAICS) code. We identified all occupational and industry codes that include livestock employment; these are listed in the supplement. Individuals whose SOCP or NAICS codes did not belong to this list were assigned a default code 00. For each combination of demographic attributes $\delta$ and employment $\epsilon$, the population is aggregated at the grid cell level. Our estimates of livestock worker counts are consistent with the American Community Survey when aggregated to the US Census Public Use Microdata Area (PUMA) level.
@@ A.2.1 Data organization and availability
[…]
@@ A.3 Gridded Livestock of the World
The Gridded Livestock of the World [24] dataset provides a gridded distribution of livestock abundance at $5$ arc minute resolution. The gridded distribution data was constructed by combining detailed livestock census statistics mined from various sources using random forest models with predictors of the following types: land use, human population, travel times, vegetation, and climate. Unsuitable areas such as water bodies and core urban centers are identified using land cover and human population density information. More details are provided in Gilbert et al. [30].
@@ 2 Constraints
3 Set Objective: Minimize $\lambda_{0}$
return $(x_{1},x_{2},\ldots,x_{m})$
@@ 3 Set Objective: Minimize $\lambda_{0}$
[…]
@@ A.5 Generation of farms
We are given farm categories $i=1,\ldots,\ell$. Let $F_{i}$ and $H_{i}$ denote the number of farms and livestock heads in category $i$. For $k=1,\ldots,\ell$, let $F_{\gamma k}$ and $H_{\gamma k}$ denote the number of farms and heads corresponding to category $k$ with respect to subtype $\gamma$ (see Table 3). The objective is to find an assignment of farms whose farm counts and composition respects these counts. Note that the categories are pairwise disjoint and cover the entire range. In addition to head counts of subtypes, we are also given counts of the livestock type as well by farm size. Since the IPF procedure does not account for these counts, there could be differences between this data and the counts obtained by aggregating farm subtype counts in our assignment. Our optimization objective is a linear combination of many parameters as discussed below.
@@ 4 For each farm $f$ in category $i$, $y_{if\gamma k}$ indicates whether $h_{if\gamma}=0$;
5 For each farm $f$ in category $i$, $z_{if\gamma k}=h_{if\gamma}$ if $h_{if\gamma}$ belongs to the category $k$ with respect to subtype $\gamma$. Otherwise, it is 00.
6 Variables for minimization objectives: $\lambda_{1}$ (alignment with known total population within each farm size category), $\lambda_{2}$ (minimize number of subtypes per farm), $\lambda_{3i}$$i\in[1,\ell]$ (equitable distribution of population in each farm size category), and (alignment with subtype population within each farm size category) $\lambda_{4}$. Constraints
@@ 7 Let $M$ be a suitably large constant;
[…]
// Subtype farm category constraints: Farm counts.
@@
8
$i\in[1,\ell],f\in[1,F_{i}],\gamma\in\Gamma$$\sum_{k}x_{if\gamma k}+y_{if\gamma k}=1$ // Farm in exactly one category
@@
// Subtype farm category constraints: Population counts
@@
// The assignment should be such that it is as close a match to the total population distribution in each category, $H_{i}$.
$i=[1,\ell],~{}\Big{|}\sum_{f,\gamma}h_{if\gamma}-H_{i}\Big{|}\leq\lambda_{1}.$ // Subtype distribution: Minimize number of subtypes per farm
@@ 9 $i=[1,\ell],~{}f=[1,F_{i}]$$\sum_{\gamma,k}x_{if\gamma k}\leq\lambda_{2}$
// Equitable distribution in each category.
@@
Set objective. Minimize $\lambda_{1}+(H+1)\cdot\lambda_{2}+(|\Gamma|+1)(H+1)\cdot\sum_{i}\lambda_{3i}+(|\Gamma|+1)(H+1)^{2}\cdot\lambda_{4}$. return $(h_{if\gamma}\mid i=[1,\ell],f=[1,F_{i}],\gamma\in\Gamma)$
@@ 5 Constraints
Set objective. Minimize $\lambda_{5}$. return $(h_{if\gamma}\mid i=[1,\ell],f=[1,F_{i}],\gamma\in\Gamma)$

### v3

@@ 2.1 The Digital Similar
The digital similar $\mathcal{DS}$ provides a unified gridded representation of livestock production and processing operations, the human population, and wild bird populations in the contiguous US. Figure 1 provides a layered view of the digital similar along with a summary of population sizes. Table 1 provides an overview of the data sources used to construct it. Formally, $\mbox{$\mathcal{DS}$}(V,\mathcal{L},\mathcal{P},\mathcal{B},\mathcal{H})$ is defined over a grid $V$ overlaid on the study region, which, in our case, is the contiguous United States. Each grid cell $v\in V$ has attributes that capture the details of each of these components. In the current setting, we use a $5\times 5$ arc minute2 grid. Descriptions of the components $\mathcal{L}$, $\mathcal{P}$, $\mathcal{B}$, and $\mathcal{H}$ are provided below.
@@
[…]
@@ 2.3 Validation of farm locations
Concentrated Agricultural Feed Operations (CAFOs) are large animal feeding operations that are a potential hazard to the environment and health. CAFOs are regulated by the Environmental Protection Agency (EPA), and some state agencies provide location information, among other attributes. We obtained such data from CafoMaps [49]. We focus on cattle, hogs, and chickens. For each county–livestock instance for which such data is present, we selected large farms from our farm assignment based on livestock-specific thresholds informed by CAFO size specifications provided by various states [52]. We computed the Haversine distance of each CAFO location to the centroid of each grid cell that contains a large farm. We construct a weighted complete bipartite graph $G(A,B)$ for each county–livestock instance. Here AA corresponds to farm locations from our assignments; each farm is assumed to be located at the centroid of the grid cell to which it belongs. The set $B$ corresponds to CAFO locations. For each $u\in A$ and $v\in B$, the weight on the edge $(u,v)$ is the inverse of the distance between the two locations. We compute a maximum weighted perfect matching [^2] For a complete bipartite graph $G(A,B,E)$, where $|A|=m$ and $|B|=n$, a perfect matching consists of $\min\{m,n\}$ edges. of this bipartite graph to match each CAFO location to a farm in $\mathcal{DS}$. The main objective is to map as many CAFO locations as possible. It is possible that the number of farms is greater than the number of CAFO locations, as not all locations are listed. The results of the matching are analyzed in Figure 5(a). We considered two sets of thresholds, the second set corresponding to larger farms compared to the first. A large percentage of CAFO locations were matched in the case of cattle (>95>95%) and chickens (>83>83%), while in the case of hogs we observe only 5050% match. A closer examination of AgCensus data reveals the reason for the low number of matches for hogs: the number of farms specified by the AgCensus dataset for the relevant size categories is less than the number of CAFO locations specified. Among the matched locations, we observe that 9090% of the CAFO locations are at most $10$ miles from the grid centroid of the corresponding farm from the digital similar, which places it in the same grid cell or a neighboring one.
@@
We note that a majority of matched CAFO farms corresponding to cattle and hogs are within 2020 miles of the matched farm in $\mathcal{DS}$.
@@ 2.4 Wild bird abundance and H5N1 incidence
The results presented in Figure 5(b) summarizes the Top-K recall results for all states and different parts of the year. We observe that, for all quarters, the larger the bird population, the greater the likelihood of observed H5N1 cases. Secondly, we observe that, in most cases, the top 10 abundant birds cover all reported H5N1 cases (for $k\geq 10$), so the recall value is close to 11. More plots are shown in the supplement for multiple US states with details about specific bird types. These results suggest a robust correlation between bird abundance and H5N1 case occurrences.
@@ 2.6 Risk Estimation
where $P(i,s)$ is the livestock population from the livestock layer $\mathcal{L}$, $A(i,t)$ is the wild bird abundance at time $t$ from the corresponding layer $\mathcal{B}$ and $\mbox{$B_{\text{H5}}$}(i,t)$ is an estimate of the proportion of the wild bird population infected with the disease at time $t$ [35, 53]. For each time period, we aggregate the risk across all grid cells of a county to obtain risk $R_{c}(s,t)$ for a county $c$. We assign risk percentile ranks to counties, designating them as “Very high” ($\geq 95$ percentile), “High” (90–94), “Medium” (75–89) and “Low” (0–74). We perform a quarterly assignment. Our results are described below. From a livestock perspective, we focus on milk cattle, turkeys and chicken layers, which have been among the most affected among the livestock subtypes.
@@
Figure 7: Results corresponding to livestock subtype-specific quarterly risk maps based on county ranks. The risk scores were ranked per quarter and counties were assigned risk profiles: “Very high” ($\geq 95\%$), “High” (909490-94), “Medium” (758975-89) and “Low”. Here, we show summary plots for milk cattle and turkeys. (a) Persistence of elevated risk: Quarter-specific ranks were combined to obtain a single ranking of counties by sorting counties by number of occurrences of “Very high” risk profiles across quarters followed by number of occurrences of “high” risk profiles and so on. (b) Period of peak risk: For each county, we plot the period for which the risk is maximum ($\operatorname*{arg\,max}_{t={1,2,3,4}}R_{c}(s,t)$). (c) Comparing ground-truth incidence reports with risk profiles. Each incidence is mapped to the risk profile of the corresponding county–quarter pair. Then, the incidences are binned into the respective risk profiles. (d) Persistence of elevated risk under a potential scenario of H5N1 spread in cattle (definition same as in (a)).
@@ 4.2 Livestock
AgCensus provides counts of heads (i.e., population size) and farms for various livestock types and subtypes. The data is available at three different administrative levels – country, state, and county. The data organization is livestock-type specific, making it a non-trivial task to extract relevant information. Farms are binned into categories based on the head counts of the corresponding livestock type. We preprocessed the data to ensure that these categories are disjoint and the categorization is identical at both state and county levels. Ideally, given a livestock subtype and administrative level, both farm and head counts are provided for each farm size category. Also provided is the total count of farms and heads. Together, we have four possible types of counts: (i) state-total: total number of farms/heads; (ii) state-by-farm-size: number of farms/heads per farm size category, and corresponding county-level counts; (iii) county-total; and (iv) county-by-farm-size. More details with examples are provided in the supplement. However, some head counts are missing in all count types. In the case of poultry, even farm size categories are missing for all subtypes except for (chicken) layers. The GLW [24] dataset provides a gridded distribution of livestock abundance at 55 arc minute resolution for the livestock types, but not for the subtypes. The details of the processing of the GLW data are provided in the supplement.
@@ 4.3 Wild Bird Abundance and Movement
We leveraged data from eBird’s Status and Trends products [27] (see eBird in Table 1) to construct this component. We recall that this component $\mathcal{B}\big(\Theta_{\mathcal{B}},A(\cdot),G_{\mathcal{B}}\big)$ captures spatiotemporal abundance and movement of multiple species of birds. The eBird data provides weekly estimates of relative bird abundance across a high-resolution grid ($\times2.96km$). These estimates, derived from ensemble machine learning models, represent the expected count of a species on a standardized eBird checklist at a given location and date. The models combine millions of citizen science observations with environmental predictors, accounting for factors such as land cover, climate variables, and observation effort. A relative abundance of 1.0 for a species at a particular location and time would indicate that an average eBird checklist at that place and time would be expected to count one individual of that species. Higher values indicate more individuals would be expected, while lower values indicate the species would be observed less frequently or in smaller numbers. This approach ensures that our model captures the most relevant species for studying avian influenza transmission. We chose bird species for which H5N1 cases were observed in the period 2022-2024. A total of 40 species of birds were identified, of which abundance data was available for 36 species. For each of the selected species, we extract relative abundance values along with their associated geographic coordinates for all 52 weeks in the year. Our processing pipeline is described in the supplement.
@@ 4.5 Human Population
We develop a gridded representation of the US population with rich demographic and employment-related attributes. This data is derived from a synthetic population [18, 32, 23] that is developed using diverse datasets such as census data, land use data, activity patterns, building maps, etc. Each individual in the population is associated with a residential location, an occupation identified by the Standard Occupational Classification code (SOCP), and an industry, identified by the North American Industry Classification System (NAICS) code. We identified all occupational and industry codes that include livestock employment; these are listed in the supplement. Individuals whose SOCP or NAICS codes did not belong to this list were assigned a default code 0. For each combination of demographic attributes $\delta$ and employment $\epsilon$, the population is aggregated at the grid cell level. Our estimates of livestock worker counts are consistent with the American Community Survey when aggregated to the US Census Public Use Microdata Area (PUMA) level.
@@ A.2.1 Data organization and availability
$i$ $W_{i}^{\mathrm{min}}$$W_{i}^{\mathrm{max}}$ $(F_{i},H_{i})$ $(F_{i,\text{beef}},H_{i,\text{beef}})$ $(F_{i,\text{milk}},H_{i,\text{milk}})$ $(F_{i,\text{other}},H_{i,\text{other}})$
@@ A.3 Gridded Livestock of the World
The Gridded Livestock of the World [24] dataset provides a gridded distribution of livestock abundance at 55 arc minute resolution. The gridded distribution data was constructed by combining detailed livestock census statistics mined from various sources using random forest models with predictors of the following types: land use, human population, travel times, vegetation, and climate. Unsuitable areas such as water bodies and core urban centers are identified using land cover and human population density information. More details are provided in Gilbert et al. [30].
@@ A.4 Filling gaps
5
61exSet Objective: Minimize $\lambda_{0}$
7
1exreturn $(x_{1},x_{2},\ldots,x_{m})$
@@
    State-total head counts. The total head count here corresponds to the country head count that is available for every livestock subtype. There are potentially four sources that can be used to derive bounds. If farm counts per farm-size category are given at the state level, an initial set of lower and upper bounds can be derived as follows: $\mbox{$W_{i}^{\mathrm{min}}$}F_{i\gamma}\leq H_{i\gamma}\leq\mbox{$W_{i}^{\mathrm{max}}$}F_{i\gamma}$. The lower bounds are refined by using the available head counts for each farm size category. Similar refinement can be done from counts per farm size at the county level. The sum of the lower bounds across farm categories provides a lower bound for the state total. Finally, if county totals are provided for some counties of the state, their sum provides another lower bound. We set the final lower bound to be the maximum of bounds obtained as above. This data is fed to FillGaps to obtain estimates for the missing counts.
@@ A.5 Generation of farms
We are given farm categories $i=1,\ldots,\ell$. Let $F_{i}$ and $H_{i}$ denote the number of farms and livestock heads in category ii. For $k=1,\ldots,\ell$, let $F_{\gamma k}$ and $H_{\gamma k}$ denote the number of farms and heads corresponding to category $k$ with respect to subtype $\gamma$ (see Table 3). The objective is to find an assignment of farms whose farm counts and composition respects these counts. Note that the categories are pairwise disjoint and cover the entire range. In addition to head counts of subtypes, we are also given counts of the livestock type as well by farm size. Since the IPF procedure does not account for these counts, there could be differences between this data and the counts obtained by aggregating farm subtype counts in our assignment. Our optimization objective is a linear combination of many parameters as discussed below.
@@ 5 For each farm $f$ in category $i$, $y_{if\gamma k}$ indicates whether $h_{if\gamma}=0$;
6 For each farm $f$ in category $i$, $z_{if\gamma k}=h_{if\gamma}$ if $h_{if\gamma}$ belongs to the category $k$ with respect to subtype $\gamma$. Otherwise, it is 0.
7 Variables for minimization objectives: $\lambda_{1}$ (alignment with known total population within each farm size category), $\lambda_{2}$ (minimize number of subtypes per farm), $\lambda_{3i}$$i\in[1,\ell]$ (equitable distribution of population in each farm size category), and (alignment with subtype population within each farm size category) $\lambda_{4}$.
81exConstraints
@@ 9 Let $M$ be a suitably large constant;
$\mbox{$W_{i}^{\mathrm{min}}$}\leq\sum_{\gamma}h_{if\gamma}\leq\mbox{$W_{i}^{\mathrm{max}}$}$ // Category farm size constraint
10
1ex// Subtype farm category constraints: Farm counts.
@@
11
[…]
@@
12
1ex// Subtype farm category constraints: Population counts
@@
13
1ex// The assignment should be such that it is as close a match to the total population distribution in each category, $H_{i}$.
14 $i=[1,\ell],~\Big|\sum_{f,\gamma}h_{if\gamma}-H_{i}\Big|\leq\lambda_{1}.$
1ex// Subtype distribution: Minimize number of subtypes per farm
@@ 15 $i=[1,\ell],~f=[1,F_{i}]$$\sum_{\gamma,k}x_{if\gamma k}\leq\lambda_{2}$
16
1ex// Equitable distribution in each category.
@@
18
191exSet objective. Minimize $\lambda_{1}+(H+1)\cdot\lambda_{2}+(|\Gamma|+1)(H+1)\cdot\sum_{i}\lambda_{3i}+(|\Gamma|+1)(H+1)^{2}\cdot\lambda_{4}$.
1exreturn $(h_{if\gamma}\mid i=[1,\ell],f=[1,F_{i}],\gamma\in\Gamma)$
@@ 4 For $1\leq i\leq N_{f}$ and $1\leq j\leq N_{c}$$h_{ij}=P_{i}$ if $x_{ij}=1$, else 0;
8
91exSet objective. Minimize $\lambda_{5}$.
1exreturn $(h_{if\gamma}\mid i=[1,\ell],f=[1,F_{i}],\gamma\in\Gamma)$